OpenStudy (anonymous):

calculate the rate of heat flow through glass window 3m x 2 m in area and 4.1 mm thick , if the temperature at the inner and outer surface are 18 degree celcius and 13 degree celcius. k for glass is 0.84 J/s.m.C......tq

OpenStudy (michele_laino):

Here we have to apply the Fourier Law, so the requested rate of heat is given by the subsequent formula: \[\Large \frac{{\Delta Q}}{{\Delta t}} = kS\frac{{{\theta _2} - {\theta _1}}}{d}\]

OpenStudy (michele_laino):

where: \[\Large \begin{gathered} {\theta _2} - \theta = 18 - 13 = ...^\circ C \hfill \\ \hfill \\ S = 3 \times 2 = ...{m^2} \hfill \\ \hfill \\ d = 4.1 \times {10^{ - 3}}\;meters \hfill \\ \hfill \\ k = 0.84\frac{{joules}}{{\sec \cdot m \cdot ^\circ C}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

thank you so much!!! @Michele_Laino